# How Should I Build A Benchmark? Revisiting Code-Related Benchmarks For LLMs

### v2

@@ Abstract
Various benchmarks have been proposed to assess the performance of large language models (LLMs) in different coding scenarios. We refer to them as code-related benchmarks. However, there are no systematic guidelines by which such a benchmark should be developed to assure its quality, reliability, and reproducibility. We propose How2Bench comprising a 55-criteria checklist as a set of guidelines to comprehensively govern the development of code-related benchmarks. Using How2Bench, we profiled 274 benchmarks released within the past decade and found concerning issues. Nearly 70% of the benchmarks did not take measures for data quality assurance; over 10% did not even open source or only partially open source. Many highly cited benchmarks have loopholes, including duplicated samples, incorrect reference codes/tests/prompts, and unremoved sensitive/confidential information. Finally, we conducted a human study involving 49 participants and revealed significant gaps in awareness of the importance of data quality, reproducibility, and transparency. For ease of use, we provide a printable version of How2Bench in Appendix E.
@@
How Should I Build A Benchmark?
Revisiting Code-Related Benchmarks For LLMs
@@ 1 Introduction
We observed that even highly cited benchmarks have loopholes, including duplicated samples, incorrect reference/tests, unclear displays, and unremoved sensitive/confidential information. We also observed these loopholes can propagate. Over 18% of the benchmarks serve as data sources (Figure 8) for subsequent benchmarks. Therefore, the data quality of benchmarks affects their credibility and likely impacts future benchmarks.
@@ 2.2 Related Studies and Surveys
Several recent surveys and empirical studies have profiled the status quo of LLM development. These studies either explore the overall performance for certain areas such as software engineering Hou et al. (2023); Wang et al. (2024a) or investigate the capabilities of LLMs on specific tasks such as code generation Dou et al. (2024); Yu et al. (2024) and test generation Schäfer et al. (2024); Yuan et al. (2024, 2023a). A survey Chang et al. (2024) about how to evaluate LLMs was proposed to answer what/where/how to evaluate LLMs. This paper differs from these studies in its purpose and perspectives. Unlike these benchmarks, our work proposed guidelines for future benchmark development and provided a checklist to assess the quality of these existing benchmarks.
@@ 3.1 The Lifecycle of Benchmark Development
Phase 0. Design. At the beginning of benchmark development, it is vital to identify the motivation, the scope and the capabilities required by the application scenario of interest.
@@ 3.2 Study Design
Our study consists of four steps (Figure 2).
@@
Step 2. Literature Profiling. This step begins by collecting related benchmarks according to their publication time, venue, and coding tasks, and then employing techniques like snowballing to ensure a comprehensive collection. This step leads to 274 code-related benchmarks for study. The detailed statistics can be found in the Appendix D. This step is followed by profiling each selected benchmark through a thorough review of corresponding papers and examination of the released artifacts or homepages associated with these benchmarks. The phase is completed by reporting statistics that highlight overall trends, advantages, and disadvantages identified during the profiling, providing a clear and structured overview of existing benchmarks.
@@
Step 4. Human Study. The final step is a human study that evaluates the importance and practicality of How2Bench. This involves designing a questionnaire to gather diverse, logical insights, which is then distributed to a targeted audience. After collecting and filtering responses for quality, the data is analyzed to derive insights. See Appendix B for details.
@@ 4.1 Guideline for Benchmark Design
Key Statistics – According to our statistics among 270+ benchmarks, apparent research bias can be observed in terms of coding tasks, programming languages, and code granularities are observed (Appendix A.1). For example, as shown in Figure 11, 36.13% (99/274) are code generation benchmarks, followed by program repair 9.85% (27/274).
@@
Lastly, Figure 12 shows that 58% (158/274) code-related benchmarks involve Python, followed by 39% (107/274) involving Java. In comparison, 31 programming languages are only covered by one benchmark, and less than five benchmarks cover other 19 programming languages. This observation consolidates the observation from previous works Cao et al. (2024a); Hou et al. (2023) on a larger scale.
@@ 4.2 Guideline for Construction
For data preprocess and cleaning, it also stresses handling code-specific aspects, such as compilability and execution, along with cleaning and manually reviewing data for quality assurance. Output validation methods and evaluation metrics must be carefully designed and reviewed to ensure they effectively measure the benchmark’s goals. Lastly, it suggests considering additional evaluation perspectives, such as safety and security checks, ensuring the code does not contain sensitive information.
@@
Also, since we focus on code-related benchmarks, which usually accompany test cases, test coverage also needs to be considered. As pointed out by prior study Liu et al. (2023a), inadequate test coverage can lead to inflated evaluation results. However, according to our statistics, we observed that only 8.7% of benchmarks have considered test coverage when using test cases as oracles. It severely affects the reliability of findings on these benchmarks, potentially misguiding future research and application decisions based on these flawed assessments.
@@
[style=MyFrame] ▲ Severity – Most benchmarks display severe loopholes in data preparation and curation processes. Essential steps such as data cleaning, process oversight, and quality checks are often neglected.
@@ 4.3 Guideline for Evaluation
Additionally, the experiment environment is essential for reproducibility and transparency. Indeed, the hardware, software, and platform environments used during experiments might influence the outcomes Ghosh (2024). Furthermore, because of the nondeterministic nature of LLMs, experiments should be repeated, and randomization strategies should be used to mitigate the effects of randomness and parameter configuration biases. Lastly, meticulously documented logs of the experimental process are advised to facilitate transparency and reproducibility, detailing everything from parameter settings to the specific LLM frameworks used.
@@
The prompt quality also greatly impacts the LLM evaluation He et al. (2024b). According to a recent study, up to 40% performance vary could be observed in code translation task He et al. (2024b). So, carefully designing a prompt needs consideration. However, 73.3% representative benchmarks (Appendix C) do not validate whether the prompts they used are well-designed (Appendix A.4). Similarly, though 94.9% benchmarks were evaluated in a zero-shot manner, only 21.2% benchmarks were evaluated under few-shot, 8.8% under Chain-of-Thought and 2.6% under RAG (Appendix A.4).
@@ 4.4 Guideline for Evaluation Analysis
Explanation – We listed 10 criteria for the evaluation analysis phase, as shown in Figure 6. Regarding the perspectives of analysis, inspired by classic measurement theory Suppes et al. (1962), we suggest four essential perspectives, including difficulty (whether a benchmark is appropriately challenging for LLMs), stability (whether the results are consistent through repeated trials), differentiability (whether benchmarks can differentiate the strengths and weaknesses of various LLMs).
@@
Key Statistics – Because experimental analysis is relatively subjective and cannot be obtained through mechanical scanning, we focus on 30 representative focus benchmarks (Appendix C), covering the highest cited and latest benchmarks in the five tasks (Section 3.2 Step 3). Figure 37 shows an example from CruxEval Gu et al. (2024) where the experimental scores can hardly be read from the figures.
@@
[style=MyFrame] ▲ Severity – The analysis of experimental data and the clarity of data presentation may receive less attention and worth consideration.
@@ 4.5 Guideline for Benchmark Release
Key Statistics – The final step involves the release of the benchmark. The fundamental requirement for releasing a benchmark is that it must be open-sourced. However, surprisingly, we observed that 5.1% of the benchmarks are only partially open-sourced (e.g., missing some subjects or tests), and 5.8% are not open-sourced at all (e.g., links/web pages are no longer active). 19.3% have not properly set up the license. Furthermore, prompts, which are necessary for reproducibility, are not disclosed in 52.6% of the benchmarks (Figure 39). Not to mention the lack of public information on experimental settings and experimental parameters (Figure 43). Furthermore, only 16.7% of the benchmarks make their logged experimental results publicly available. See Appendix A.6 for details.
@@ A.3.1 Data Preprocessing
Data Contamination Resolution. Data contamination Golchin and Surdeanu (2023); Cao et al. (2024b) threat has been widely discussed. A benchmark with contaminated data may yield overclaimed results, misleading the understanding of the LLMs’ capabilities. According to our statistics (Figure 23 on benchmarks from the year 2023 to 2024 (the duration when most LLMs were launched), most (78 %) benchmarks were not aware of and have not taken any measures to alleviate data contamination, being vulnerable to data contamination threat.
@@ A.4 Statistics about Evaluation
Studied LLMs. We summarize the number of LLMs that have been evaluated in each benchmark evaluation. As shown in Figure 29, the majority of benchmarks are tested with a smaller number of LLMs, as evidenced by the bars for 1 to 9 LLMs. Interestingly, there is a significant peak at the 10-49 LLM range, with 56 benchmarks falling into this category, suggesting a common practice of testing benchmarks extensively but not excessively. The drop in benchmarks tested with 50-99 and 100+ LLMs indicates that very extensive testing across numerous models is less common, potentially due to resource constraints or diminishing returns on insights gained from additional models.
@@ B.3 Interview Process
Survey Screening Since the requirement was for participants who have published papers, responses from those selecting "No" to having published a paper were excluded. Additionally, incomplete surveys where not all questions were answered were also considered invalid and excluded from the analysis.
@@ B.4 Interview Result Analysis
In total, we collected 50 responses. The respondents were from seven regions, including the United States, the United Kingdom, Germany, Australia, China, and others, as shown in Figure 49. Only one survey was invalid due to the respondent selecting “have not published a paper”, leaving 49 valid surveys for analysis. A breakdown of the respondents’ demographics is shown in Figure 50. And the detailed response for all 55 criteria in How2Bench are shown in Figure 51 and Figure 52.

### v3

# How Should We Build A Benchmark? Revisiting 274 Code-Related Benchmarks For LLMs

@@ Abstract
Various benchmarks have been proposed to assess the performance of large language models (LLMs) in different coding scenarios. We refer to them as code-related benchmarks. However, there are no systematic guidelines by which such a benchmark should be developed to assure its quality, reliability, and reproducibility. We propose How2Bench comprising a 55-criteria checklist as a set of guidelines to comprehensively govern the development of code-related benchmarks. Using How2Bench, we profiled 274 code-related benchmarks released within the past decade and found concerning issues. Nearly 70% of the benchmarks did not take measures for data quality assurance; over 10% did not even open source or only partially open source. Many highly cited benchmarks have loopholes, including duplicated samples, incorrect reference codes/tests/prompts, and unremoved sensitive/confidential information. Finally, we conducted a human study involving 49 participants and revealed significant gaps in awareness of the importance of data quality, reproducibility, and transparency. For ease of use, we provide a printable version of How2Bench in Appendix E.
@@
How Should We Build A Benchmark?
Revisiting 274 Code-Related Benchmarks For LLMs
@@ 1 Introduction
We observed that even highly cited benchmarks have loopholes, including duplicated samples, incorrect reference/tests, unclear displays, and unremoved sensitive/confidential information. We also observed these loopholes can propagate. Over 18% of the benchmarks serve as data sources for subsequent benchmarks (Figure 8). Therefore, the data quality of benchmarks affects their credibility and likely impacts future benchmarks.
@@ 2.2 Related Studies and Surveys
Several recent surveys and empirical studies have profiled the status quo of LLM development. These studies either explore the overall performance for certain areas such as software engineering Hou et al. (2023); Wang et al. (2024a) or investigate the capabilities of LLMs on specific tasks such as code generation Dou et al. (2024); Yu et al. (2024) and test generation Schäfer et al. (2024); Yuan et al. (2024b, 2023a). A survey Chang et al. (2024) about how to evaluate LLMs was proposed to answer what/where/how to evaluate LLMs. This paper differs from these studies in its purpose and perspectives. Unlike these benchmarks, our work proposed guidelines for future benchmark development and provided a checklist to assess the quality of these existing benchmarks.
@@ 3.1 The Lifecycle of Benchmark Development
Phase 0. Design. At the beginning of benchmark development, it is vital to identify the motivation, the scope and the capabilities required by the application scenario of interest. To achieve this objective, one needs to carefully consider the application scenarios, making sure these scenarios align with real-world demands. Also, it is also necessary to assess whether other benchmarks already exist that address similar tasks, and to identify any shortcomings they may possess. Furthermore, this new benchmark should be designed to evaluate specific LLMs’ capabilities; the crafted tasks are expected to reflect these capabilities.
@@ 3.2 Study Design
Our study consists of four steps (Figure 2). All steps are explained as follows.
@@
Step 2. Literature Profiling. This step begins by collecting related benchmarks according to their publication time, venue, and coding tasks, and then employing techniques like snowballing to ensure a comprehensive collection. This step leads to 274 code-related benchmarks for study. The detailed statistics can be found in the Appendix D. This step is followed by profiling each selected benchmark through a thorough review of corresponding papers and examination of the released artifacts or homepages associated with these benchmarks. The phase is completed by reporting statistics that highlight overall trends, pros, and cons identified during the profiling, providing a structured overview of existing benchmarks.
@@
Step 4. Human Study. The final step is a human study that evaluates the importance and practicality of How2Bench. This involves designing a questionnaire by first initiating and iterating to gather diverse, logical insights, which is then distributed to a targeted audience. After collecting and filtering responses for quality, the data is analyzed to derive insights. See Appendix B for details.
@@ 4.1 Guideline for Benchmark Design
Key Statistics – According to our statistics among 270+ benchmarks, apparent research bias can be observed in terms of coding tasks, programming languages, and code granularities are observed (Appendix A.1). For example, 36.13% (99/274) are code generation benchmarks, followed by program repair, with 9.85% (27/274).
@@
Lastly, Figure 12 shows that 58% (158/274) code-related benchmarks involve Python, followed by 39% (107/274) involving Java. Yet, 31 programming languages are only covered by one benchmark, and less than five benchmarks cover other 19 programming languages. This observation consolidates the observation from previous works Cao et al. (2024a); Hou et al. (2023) on a larger scale.
@@ 4.2 Guideline for Construction
For data preprocess and cleaning, it also stresses handling code-specific aspects, such as compilability and execution, along with cleaning and manually reviewing data for quality assurance. Output validation methods and evaluation metrics must be carefully designed and reviewed to ensure they effectively measure the benchmark’s goals. Lastly, it suggests considering additional evaluation perspectives, such as safety Wei et al. (2024); Yuan et al. (2024a) checks, ensuring the code does not contain sensitive information.
@@
Also, since we focus on code-related benchmarks, which usually accompany test cases, test coverage also needs to be considered. As pointed out by prior study Liu et al. (2023a), inadequate test coverage can lead to inflated evaluation results. However, we observed that only 8.7% of benchmarks have considered test coverage when using test cases as oracles (Appendix A.3). It severely affects the reliability of findings on these benchmarks, potentially misguiding future research and applications based on these flawed assessments.
@@
[style=MyFrame] ▲ Severity – Most benchmarks display severe loopholes in data preparation and curation. Quality checks are often neglected.
@@ 4.3 Guideline for Evaluation
Additionally, the experiment environment is essential for reproducibility and transparency. Indeed, the hardware, software, and platform environments used during experiments might influence the outcomes Ghosh (2024). Furthermore, because of the nondeterministic nature of LLMs, experiments should be repeated, and randomization strategies should be used to mitigate the effects of randomness and parameter configuration biases. Lastly, meticulously documented logs of the experimental process are advised to facilitate transparency and reproducibility, detailing everything from parameter settings to the specific LLM pipelines such as vLLM (Kwon et al., 2023) used.
@@
The prompt quality also greatly impacts the LLM evaluation He et al. (2024b). According to a recent study, up to 40% performance vary could be observed in code translation task He et al. (2024b). So, carefully designing a prompt needs consideration. However, 73.3% representative benchmarks (Appendix C) do not validate whether the prompts they used are well-designed (Appendix A.4). Similarly, though 94.9% benchmarks were evaluated in a zero-shot manner, only 21.2% benchmarks were evaluated under few-shot, 8.8% under Chain-of-Thought and 2.6% under RAG (Appendix A.4). However, as shown in Figure 34, 73.3% representative benchmarks (Appendix C) do not validate whether the prompt they used is well-designed.
@@ 4.4 Guideline for Evaluation Analysis
Explanation – The analysis of the experiment results is expected to be objective and comprehensive, hopefully providing insights or actionable advice. So, we listed 10 criteria for the evaluation analysis phase, as shown in Figure 6. Regarding the perspectives of analysis, inspired by classic measurement theory Suppes et al. (1962), we suggest four essential perspectives, including difficulty (whether a benchmark is appropriately challenging for LLMs), stability (whether the results are consistent through repeated trials), differentiability (whether benchmarks can differentiate the strengths and weaknesses of various LLMs), and inspiration (e.g., the correlations between the upper-/down-stream coding tasks and LLM scores).
@@
Key Statistics – Because experimental analysis is relatively subjective and cannot be obtained through mechanical scanning, we focus on 30 representative focus benchmarks (Appendix C), covering the highest cited and latest benchmarks in top five tasks. Figure 37 shows an example from CruxEval Gu et al. (2024) where the experimental scores can hardly be read from the figures.
@@
[style=MyFrame] ▲ Severity – The analysis of experimental data and the clarity of data presentation may receive less attention and worth consideration. Even in papers cited 1k+ times like MBPP Austin et al. (2021), there are instances of unclear evaluation analysis and display.
@@ 4.5 Guideline for Benchmark Release
Key Statistics – The final step involves the release of the benchmark. The fundamental requirement for releasing a benchmark is that it must be open-sourced. However, surprisingly, we observed that 5.1% of the benchmarks are only partially open-sourced (e.g., missing some subjects or tests), and 5.8% are not open-sourced at all (e.g., links/web pages are no longer active). 19.3% have not properly set up the license. Furthermore, prompts, which are necessary for reproducibility, are not disclosed in 52.6% of the benchmarks (Figure 39). Not to mention the lack of public information on experimental settings (Figure 32 and Figure 31) and experimental parameters (Figure 43). What is worse, 19.3% benchmarks do not setup licenses (Figure 44). The absence of licensing may lead to severe legal and ethical issues, potentially resulting in unauthorized use and distribution of proprietary technologies. Additionally, only 16.7% of the benchmarks make their logged experimental results publicly available (Appendix A.6).
@@ A.3.1 Data Preprocessing
Data Contamination Resolution. Data contamination Golchin and Surdeanu (2023); Cao et al. (2024b) threat has been widely discussed. A benchmark with contaminated data may yield overclaimed results, misleading the understanding of the LLMs’ capabilities. According to our statistics (Figure 23 on benchmarks from the year 2023 to 2024 (the duration when most LLMs were launched), most (81.8 %) benchmarks were not aware of and have not taken any measures to alleviate data contamination, being vulnerable to data contamination threat.
0
@@ A.4 Statistics about Evaluation
Studied LLMs. We summarize the number of LLMs that have been evaluated in each benchmark evaluation. Among the 274 benchmarks, 183 of them are evaluated over LLMs, so we show the statistics over them. As shown in Figure 29, over 34% of the benchmarks were evaluated on fewer than 3 LLMs, with 11.48% benchmarks only evaluated on one LLM. Such evaluation results can hardly be generalized to other LLMs. Furthermore, more than half of the benchmarks studied fewer than 6 LLMs (51% = (21 +22 + 20 + 4 + 12+15)/183).
@@ B.3 Interview Process
Survey Screening Since the requirement was for participants who have published papers, responses from those selecting “No” to having published a paper were excluded. Also, incomplete surveys where not all questions were answered were also considered invalid and excluded from the analysis.
@@ B.4 Interview Result Analysis
In total, we collected 50 responses. The respondents were from seven regions, including the United States, the United Kingdom, Germany, Australia, China, and others, as shown in Figure 49. Only one survey was invalid due to the respondent selecting “have not published a paper”, leaving 49 valid surveys for analysis. A breakdown of the respondents’ demographics is shown in Figure 50. The detailed responses for all 55 criteria in How2Bench are shown in Figure 51 and Figure 52.